OpenStudy (anonymous):

Joy has been saving her earnings to make a down payment on a new car. Which loan terms will result in highest total cost for the loan?

OpenStudy (rane):

Okay. This is what I would do. For A. 2.5 percent loan for 24 months. 2.5 / 24 = 0.1041666, so A is a 0.1041666 percent loan er month. Now B, which is a 3.5 percent loan for 36 months. I do the same thing. 3.5 / 36. This would be 0.0972222, which is less than A, so that would mean B is not the highest, so keep A and forget about B. Now for C. 4.5 / 48 = 0.09375. That is also less than A, so so far it seems that A is winning the compeitition. Now for D. 5.5 / 60 = 0.0916666 per month. This is less than A too, so the answer is A. The answer is A. 2.5 percent loan for 24 months.

OpenStudy (mexhitokiri):

D 5.5 percent loan for 60 months is correct.

OpenStudy (anonymous):

it is D just took this quiz